### v1

@@ 10 Final Considerations
This paper presents a practical interactive approach, whose implementation is available at https://github.com/unibz-core/CM-Mining, for automating the empirical discovery of modeling patterns in conceptual models by combining state-of-the-art graph manipulation techniques and frequent subgraph mining. Structural patterns are reusable representations of bad or good practices. They are typically harvested by manual labor-intensive processes, which usually take a long time to converge into pattern catalogs. Our automated process for pattern discovery facilitates the construction of pattern catalogs tailored for modeling languages. Additionally, we create a mechanism for helping language designers to create higher-granularity primitives in that language, i.e., modeling patterns that can become part of the grammar and tools of that language [7].
Based on the encouraging results presented here, we created a list of tasks for immediate future work. First, we are going to further assess the approach to finding unexpected patterns. In order to address this, we will involve representative users and run an evaluation through usability testing. Second, we are going to test the approach with models encoded in different conceptual modeling languages, such as BPMN or ArchiMate (R5). In the current work, we used a specific subset of (Onto)UML as a target use case scenario, leveraging a large dataset of high-quality data. However, we have designed our solution to rely on graph-structured data, enabling the encoding of any conceptual model construct. Moreover, the whole approach, which can be adapted to different input formats, is completely language-independent except for the first importing step.

### v2

@@ 10 Final Considerations
This paper presents a practical interactive approach, whose implementation is available at https://github.com/unibz-core/CM-Mining, for automating the empirical discovery of recurrent structures in conceptual models by combining state-of-the-art graph manipulation techniques and frequent subgraph mining. Structural patterns can be exploited to identify reusable representations of bad or good modeling practices. They are typically harvested by manual labor-intensive processes, which usually take a long time to converge into pattern catalogs. Our automated process for pattern discovery facilitates the construction of pattern catalogs tailored for modeling languages. Additionally, we create a mechanism for helping language engineers to create higher-granularity primitives in that language, i.e., modeling patterns that can become part of the grammar and tools of that language [7].
Based on the encouraging results presented here, we created a list of tasks for immediate future work. First, we are going to further assess the approach to finding unexpected patterns. In order to address this, we will involve representative users and run an evaluation through usability testing. Second, we are going to test the approach with models encoded in new conceptual modeling languages, such as BPMN (R5).
We chose to validate the proposed approach using two state-of-the-art conceptual modeling languages, OntoUML and ArchiMate. However, our solution is designed around graph-structured data, allowing it to encode any conceptual model constructs. Furthermore, the approach is fully language-independent, apart from the initial importing and the final visualization steps, and can be adapted to various input formats.
Our focus on these two languages and their associated datasets was driven by several key factors. First, both languages include constructs—such as classes, annotations, and cardinalities—that are also prevalent in widely-used languages like UML and BPMN. Additionally, these languages are accompanied by two high-quality, scientifically validated datasets that adhere to FAIR principles. Finally, we had direct access to the language designers, who provided essential support and feedback on the functionalities and output structures required throughout our research.